A 0.1216g sample of a monoprotic acid is dissolved in 25mL of water and the resulting solution is titrated with 0.1104M NaOH solution. A 12.5mL volume of the base is required to neutralise the acid. Calculate the molar mass of the acid.
1
Expert's answer
2020-08-17T13:34:55-0400

Monoprotic Acid Is React With Strong Base Form Salt And Water .


1 mole of Acid React with 1 mol of Base


Moles Of Acid = 0.1216 / M ....................... 1 .


Where M is Molar Mass


Moles Of Base = 12.5 * 0.1104 / 1000 . ----------------2 .


eqn 1 = eqn 2 .


0.1216 / M = 12.5 * 0.1104 / 1000 .


M = 88.15 g / mol .